# Incorporating Arbitrary Matrix Group Equivariance into KANs

### v3

@@ 6.3 Top quark tagging
The research on top quark tagging (Kasieczka et al., 2019) involves classifying hadronic tops from the QCD background. In particle collision experiments, top quark decays or other events produce several jet constituents. We observe the four-momenta $p_{1}^{\mu},p_{2}^{\mu},p_{3}^{\mu}\in\mathbb{R}^{4}$ of the three jet constituents with the highest transverse momentum $p_{T}$, and predict the event label ($1$ for top, 00 for QCD). The category of the event will not change when all jet constituents undergo the same Lorentz transformation. Consequently, this task possesses $\mathrm{O}(1,3)$ invariance (see Section A.2 for more details), with the feature space $U_{i}=3T_{1}$ and the label space $U_{o}=T_{0}$.
Similar to particle scattering, we embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN and EMLP. Furthermore, we embed $\mathrm{O}(1,3)$ equivariance into CGENN (Ruhe et al., 2023). We sample training sets of different sizes from the original dataset for evaluation. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. We set the width of the middle layer in EKAN to $n_{1}=200$ (shape as $[12,200,1]$) and the width of the middle layer in KAN to $n_{1}=384$ (shape as $[12,384,1]$) to control the number of parameters. Both MLP, EMLP and CGENN have the depth of $L=4$ and the middle layer width $n_{1}=n_{2}=n_{3}=200$ (shape as $[12,200,200,200,1]$). We apply the sigmoid function to the model’s output and use BCE as the loss function for binary classification. More implementation details are provided in Section E.3.
We report the mean $\pm$ std of the test accuracy over three runs with different random seeds, as well as the number of parameters of the models in Table 3. Since we have not observed all the jet constituents, the relationship between the labels and input features cannot be accurately expressed as an explicit function. In this case of non-symbolic formula representation, KAN cannot achieve higher accuracy with fewer parameters than MLP as expected. On the other hand, our EKAN achieves comparable results with EMLP and CGENN using fewer than $40\%$ of the parameters, improving test accuracy by $0.23\%\sim 6.44\%$ on small datasets (training set size $<10^{4}$), while decreasing by $0.19\%$ on large datasets (training set size $=10^{4}$).
@@ 7 Conclusion
To our knowledge, this work is the first attempt to combine equivariance and KANs. We view the KAN layer as a combination of spline functions and linear weights, and accordingly define the (gated) input space, post-activation space, and (gated) output space of the EKAN layer. Gated basis functions ensure the equivariance between the gated input space and the post-activation space, while equivariant linear weights guarantee the equivariance between the post-activation space and the gated output space. The prior work has demonstrated that “EMLP $>$ MLP” on tasks with symmetries and “KAN $>$ MLP” on symbolic formula representation tasks. Our experimental results further indicate that on symbolic formula representation tasks with symmetries, “EKAN $>$ EMLP” and “EKAN $>$ KAN”. Moreover, on non-symbolic formula representation tasks with symmetries, although it may be that “KAN $<$ MLP”, we show that “EKAN $>$ EMLP”. We expect that EKAN can become a general framework for applying KANs to more fields, such as computer vision and natural language processing, just as EMLP unifies classic works like CNNs and DeepSets.
@@ A.2 Groups $\mathrm{SO}^{+}(1,3)$, $\mathrm{SO}(1,3)$, and $\mathrm{O}(1,3)$
$\begin{cases}A_{1}=\begin{bmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},A_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\end{bmatrix},A_{3}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\end{bmatrix},\\ A_{4}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\\ 0&-1&0&0\\ 0&0&0&0\end{bmatrix},A_{5}=\begin{bmatrix}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&-1&0&0\end{bmatrix},A_{6}=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{bmatrix},\end{cases}$ (24)
@@
Appendix B Equivariant Multi-Layer Perceptrons (EMLP)
@@
Appendix C Proof of Theorem 4.1
@@
Appendix D Grid update
@@
Appendix E Implementation details
E.1 Particle scattering
@@
In particle scattering, we generate training sets of different sizes, and the corresponding test sets have the same sizes as the training sets. We train EKAN using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $<1000$, we set the number of epochs to $7000$, while for datasets with the training set size $\geq 1000$, we set the number of epochs to $15000$, which is sufficient for the MSE loss to converge to the minimum. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.
E.2 Three-body problem
@@
E.3 Top quark tagging
In top quark tagging, we train all models using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $\leq 1000$, we set the number of epochs to $1000$, while for datasets with the training set size $>1000$, we set the number of epochs to $2000$, which is sufficient for the BCE loss to converge to the minimum. The grids of EKAN and KAN are updated every $5$ epochs and stop updating at the $50$th epoch. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.

### v4

@@ 6.3 Top Quark Tagging
The research on top quark tagging (Kasieczka et al., 2019) involves classifying hadronic tops from the QCD background. In particle collision experiments, top quark decays or other events produce several jet constituents. We observe the four-momenta $p_{1}^{\mu},p_{2}^{\mu},p_{3}^{\mu}\in\mathbb{R}^{4}$ of the three jet constituents with the highest transverse momentum $p_{T}$, and predict the event label ($1$ for top, 0 for QCD). The category of the event will not change when all jet constituents undergo the same Lorentz transformation. Consequently, this task possesses $\mathrm{O}(1,3)$ invariance (see Section A.2 for more details), with the feature space $U_{i}=3T_{1}$ and the label space $U_{o}=T_{0}$. The top quark tagging dataset is sourced from Kasieczka et al. (2019), and we assume that only the $n_{comp}=3$ jet constituents with the highest transverse momentum $p_{T}$ are observed, which differs from the setup in CGENN (Ruhe et al., 2023) and LorentzNet (Gong et al., 2022), where all $n_{comp}=200$ jet constituents are available.
Similar to particle scattering, we embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN and EMLP. Furthermore, we embed $\mathrm{O}(1,3)$ equivariance into CGENN (Ruhe et al., 2023). We sample training sets of different sizes from the original dataset for evaluation. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. We set the width of the middle layer in EKAN to $n_{1}=200$ (shape as $[12,200,1]$) and the width of the middle layer in KAN to $n_{1}=384$ (shape as $[12,384,1]$) to control the number of parameters. Both MLP, EMLP and CGENN have the depth of $L=4$ and the middle layer width $n_{1}=n_{2}=n_{3}=200$ (shape as $[12,200,200,200,1]$). We apply the sigmoid function to the model’s output and use BCE as the loss function for binary classification. More implementation details are provided in Section G.3.
We report the mean $\pm$ std of the test accuracy over three runs with different random seeds, as well as the number of parameters of the models in Table 3. Since we have not observed all the jet constituents, the relationship between the labels and input features cannot be accurately expressed as an explicit function. In this case of non-symbolic formula representation, KAN cannot achieve higher accuracy with fewer parameters than MLP as expected. On the other hand, our EKAN achieves comparable results with EMLP and CGENN using fewer than $40\%$ of the parameters, improving test accuracy by $0.23\%\sim 6.44\%$ on small datasets (training set size $10^{4}$), while decreasing by $0.19\%$ on large datasets (training set size $10^{4}$).
@@ 7 Conclusion
To our knowledge, this work is the first attempt to combine equivariance and KANs. We view the KAN layer as a combination of spline functions and linear weights, and accordingly define the (gated) input space, post-activation space, and (gated) output space of the EKAN layer. Gated basis functions ensure the equivariance between the gated input space and the post-activation space, while equivariant linear weights guarantee the equivariance between the post-activation space and the gated output space. The prior work has demonstrated that “EMLP >> MLP” on tasks with symmetries and “KAN >> MLP” on symbolic formula representation tasks. Our experimental results further indicate that on symbolic formula representation tasks with symmetries, “EKAN >> EMLP” and “EKAN >> KAN”. Moreover, on non-symbolic formula representation tasks with symmetries, although it may be that “KAN << MLP”, we show that “EKAN >> EMLP”. We expect that EKAN can become a general framework for applying KANs to more fields, such as computer vision and natural language processing, just as EMLP unifies classic works like CNNs and DeepSets.
8 Limitations
One limitation of our approach is that the use of gating mechanisms can reduce the expressive power of EKAN, as discussed in Appendix D of EMLP (Finzi et al., 2021). However, due to the inherently complex structure of KANs (which involve intricate B-spline formulations), introducing symmetries into KANs is challenging. Therefore, for the sake of simplicity and clarity, we opt to incorporate gated non-linearities in a hierarchical manner. We anticipate that future improvements could enhance the expressive power and flexibility of EKAN.
Another limitation is that when data is sufficient, the advantages of EKAN diminish compared with baselines. We provide additional experimental results in Appendix H to illustrate this point. This implies that EKAN enhances generalization on small data but does not unlock the upper limit of expressive power.
Acknowledgements
Z. Lin was supported by National Key R&D Program of China (2022ZD0160300) and the NSF China (No. 62276004).
@@ A.2 Groups $\mathrm{SO}^{+}(1,3)$, $\mathrm{SO}(1,3)$, and $\mathrm{O}(1,3)$
[…]
@@
Appendix B Concrete Examples of Space Structure
First, let’s intuitively understand the dual (*), direct sum ($\oplus$), and tensor product ($\otimes$) operations. Consider two vector spaces $X=\mathbb{R}^{2},Y=\mathbb{R}^{3}$, and vectors $x=(x_{1},x_{2})\in X,y=(y_{1},y_{2},y_{3})\in Y$. Then, all $x\oplus y=(x_{1},x_{2},y_{1},y_{2},y_{3})$ form the space $X\oplus Y=\mathbb{R}^{5}$, all $x\otimes y=(x_{1}y,x_{2}y)=(x_{1}y_{1},x_{1}y_{2},x_{1}y_{3},x_{2}y_{1},x_{2}y_{2},x_{2}y_{3})$ form the space $X\otimes Y=\mathbb{R}^{6}$, and all the coefficients $\mathrm{vec}(W)$ of the linear maps $Wx=y$ form the space $Y\otimes X^{*}=\mathbb{R}^{6}$.
Then, if we define how the group transformation acts on $X,Y$, the form of its action on these composite spaces can naturally be derived. Equation 4 provides the derivation rules. This paper assumes the group to be a matrix group. If a group element $g\in G$ acts on a vector $x\in X$ in the form of its corresponding linear transformation $gx$, then we call $X$ the base vector space of $G$.
We have defined the “addition” and “multiplication” between spaces. Thus, for the base vector space $V$ of a group $G$, any complex space structure can be organized into the form of a “polynomial” with respect to $V$, which is the origin of Equation 3. Note that Equation 3 simultaneously defines $T(p,q)=V^{p}\otimes(V^{*})^{q}$. We abbreviate $T(p,0)$ as $T_{p}$.
In the vast majority of scenarios, the feature space of a dataset takes simple forms such as vector stacking $cT_{1}=cV$ or matrices $T_{2}=V\otimes V$, while complex spaces like $T(p,q)$ rarely appear. However, the latent spaces between equivariant layers can be highly intricate (e.g., they may have $384$ dimensions), and their decomposition forms with respect to the base vector space $V$ often involve $T(p,q)$ (the decomposition is automatically handled by the software based on dimensionality, as described in Section 6). The practical implication is that, according to the rules of Equation 4, we define how group transformations operate on the latent spaces.
It is not correct that any (continuous, real, finite-dimensional) representation $U$ of a matrix group can be written as in Equation 3. It can be shown (in the case of a compact Lie group for example) that $U$ is a subrepresentation of the direct sum on the right hand side. However, it is worth noting that, for the vast majority of practical applications, considering input/output representations of this form should suffice.
Appendix C Equivariant Multi-Layer Perceptrons (EMLP)
@@
Appendix D Proof of Theorem 4.1
@@
Appendix E Grid Update
@@
Appendix F Equivariant Error Evaluation
For all trained models $f_{\theta}$ in Section 6, we use $\mathcal{L}_{equi}=\mathbb{E}_{x,g}\|\rho_{o}(g)f_{\theta}(x)-f_{\theta}(\rho_{i}(g)x)\|^{2}$ to evaluate their equivariant errors. The experimental results are presented in Tables 4, 5 and 6, which indicate that our EKAN and EMLP can structurally guarantee strict equivariance, whereas non-equivariant models cannot. For particle scattering, the results of EMLP and MLP are sourced from the original paper (Finzi et al., 2021), so we do not present their equivariant errors here.
Table 4: Equivariant error of different models on the particle scattering dataset with different training set sizes. We present the results in the format of mean $\pm$ std.
Models Training set size
$10^{2}$ $10^{2.5}$ $10^{3}$ $10^{3.5}$ $10^{4}$
KAN $(3.14\pm 0.08)\times 10^{-1}$ $1.02\pm 0.69$ $(4.88\pm 2.02)\times 10^{-1}$ $(6.90\pm 3.25)\times 10^{-1}$ $(1.64\pm 0.41)\times 10^{-1}$
KAN + augmentation $(2.78\pm 0.30)\times 10^{-1}$ $(8.81\pm 8.41)\times 10^{-1}$ $(1.08\pm 0.55)\times 10^{-1}$ $(1.32\pm 0.52)\times 10^{-1}$ $(2.36\pm 0.17)\times 10^{-1}$
EKAN-$\mathrm{SO}^{+}(1,3)$ (Ours) $(6.13\pm 1.76)\times 10^{-14}$ $(9.32\pm 2.75)\times 10^{-14}$ $(7.48\pm 1.61)\times 10^{-14}$ $(8.33\pm 0.69)\times 10^{-14}$ $(6.57\pm 0.06)\times 10^{-14}$
EKAN-$\mathrm{SO}(1,3)$ (Ours) $(1.13\pm 0.78)\times 10^{-13}$ $(7.64\pm 1.67)\times 10^{-14}$ $(6.86\pm 0.11)\times 10^{-14}$ $(7.98\pm 0.45)\times 10^{-14}$ $(6.62\pm 0.89)\times 10^{-14}$
EKAN-$\mathrm{O}(1,3)$ (Ours) $(5.67\pm 0.87)\times 10^{-14}$ $(7.95\pm 2.67)\times 10^{-14}$ $(5.88\pm 0.81)\times 10^{-14}$ $(9.00\pm 0.63)\times 10^{-14}$ $(8.43\pm 0.18)\times 10^{-14}$
Table 5: Equivariant error of different models with different numbers of parameters on the three-body problem dataset. We present the results in the format of mean $\pm$ std.
Models Number of parameters
$10^{4.5}$ $10^{4.75}$ $10^{5}$ $10^{5.25}$ $10^{5.5}$
MLP $(1.44\pm 0.13)\times 10^{-3}$ $(1.51\pm 0.14)\times 10^{-3}$ $(1.54\pm 0.03)\times 10^{-3}$ $(1.33\pm 0.17)\times 10^{-3}$ $(1.33\pm 0.10)\times 10^{-3}$
MLP + augmentation $(3.16\pm 0.32)\times 10^{-3}$ $(3.47\pm 0.24)\times 10^{-3}$ $(3.43\pm 0.24)\times 10^{-3}$ $(3.35\pm 0.16)\times 10^{-3}$ $(3.32\pm 0.26)\times 10^{-3}$
EMLP-$\mathrm{SO}(2)$ $(2.80\pm 2.18)\times 10^{-13}$ $(1.57\pm 1.39)\times 10^{-13}$ $(2.17\pm 0.37)\times 10^{-14}$ $(2.99\pm 2.51)\times 10^{-14}$ $(1.87\pm 0.61)\times 10^{-14}$
EMLP-$\mathrm{O}(2)$ $(7.87\pm 6.26)\times 10^{-13}$ $(1.87\pm 1.32)\times 10^{-12}$ $(3.22\pm 4.53)\times 10^{-11}$ $(1.46\pm 2.03)\times 10^{-12}$ $(7.23\pm 6.13)\times 10^{-14}$
KAN $(3.00\pm 2.28)\times 10^{-1}$ $(1.21\pm 0.26)\times 10^{-2}$ $(5.90\pm 0.78)\times 10^{-3}$ $(5.93\pm 1.45)\times 10^{-3}$ $(4.03\pm 0.81)\times 10^{-3}$
KAN + augmentation $(2.78\pm 0.10)\times 10^{-3}$ $(2.49\pm 0.11)\times 10^{-3}$ $(2.37\pm 0.11)\times 10^{-3}$ $(2.35\pm 0.11)\times 10^{-3}$ $(2.29\pm 0.08)\times 10^{-3}$
EKAN-$\mathrm{SO}(2)$ (Ours) $(3.80\pm 3.84)\times 10^{-13}$ $(3.02\pm 3.05)\times 10^{-13}$ $(9.66\pm 5.85)\times 10^{-14}$ $(3.33\pm 1.47)\times 10^{-14}$ $(3.31\pm 1.31)\times 10^{-14}$
EKAN-$\mathrm{O}(2)$ (Ours) $(9.79\pm 6.13)\times 10^{-13}$ $(1.46\pm 1.30)\times 10^{-13}$ $(2.99\pm 2.77)\times 10^{-13}$ $(1.75\pm 1.55)\times 10^{-13}$ $(7.62\pm 4.91)\times 10^{-14}$
Table 6: Equivariant error of different models on the top quark tagging dataset ($n_{comp}=3$) with different training set sizes. We present the results in the format of mean $\pm$ std.
Models Training set size
$10^{2}$ $10^{2.5}$ $10^{3}$ $10^{3.5}$ $10^{4}$
MLP $(4.94\pm 0.40)\times 10^{-1}$ $(4.00\pm 0.17)\times 10^{-1}$ $(3.73\pm 0.05)\times 10^{-1}$ $(3.24\pm 0.11)\times 10^{-1}$ $(1.87\pm 0.03)\times 10^{-1}$
MLP + augmentation $(4.73\pm 0.28)\times 10^{-1}$ $(2.15\pm 0.12)\times 10^{-1}$ $(1.90\pm 0.55)\times 10^{-1}$ $(3.73\pm 0.39)\times 10^{-1}$ $(3.40\pm 0.63)\times 10^{-1}$
EMLP-$\mathrm{SO}^{+}(1,3)$ $(2.92\pm 2.30)\times 10^{-6}$ $(1.09\pm 0.84)\times 10^{-7}$ $(3.88\pm 1.44)\times 10^{-9}$ $(3.54\pm 2.47)\times 10^{-9}$ $(1.65\pm 2.11)\times 10^{-8}$
EMLP-$\mathrm{SO}(1,3)$ $(6.81\pm 8.20)\times 10^{-7}$ $(1.36\pm 1.82)\times 10^{-7}$ $(5.49\pm 2.92)\times 10^{-9}$ $(2.69\pm 1.48)\times 10^{-9}$ $(1.71\pm 0.36)\times 10^{-9}$
EMLP-$\mathrm{O}(1,3)$ $(2.41\pm 1.68)\times 10^{-7}$ $(1.86\pm 1.80)\times 10^{-7}$ $(6.20\pm 0.59)\times 10^{-9}$ $(1.14\pm 1.36)\times 10^{-8}$ $(1.26\pm 0.28)\times 10^{-9}$
KAN $(1.36\pm 1.78)\times 10^{-1}$ $(1.35\pm 1.77)\times 10^{-1}$ $(1.34\pm 1.77)\times 10^{-1}$ $(9.90\pm 14.00)\times 10^{-5}$ $(1.27\pm 1.79)\times 10^{-1}$
KAN + augmentation $(1.20\pm 1.61)\times 10^{-1}$ $(2.10\pm 2.97)\times 10^{-3}$ $(1.10\pm 1.51)\times 10^{-1}$ $(1.52\pm 2.15)\times 10^{-4}$ $(1.44\pm 1.78)\times 10^{-1}$
EKAN-$\mathrm{SO}^{+}(1,3)$ (Ours) $(3.41\pm 2.99)\times 10^{-7}$ $(1.66\pm 0.85)\times 10^{-8}$ $(1.62\pm 0.51)\times 10^{-9}$ $(1.65\pm 1.01)\times 10^{-9}$ $(1.64\pm 1.53)\times 10^{-9}$
EKAN-$\mathrm{SO}(1,3)$ (Ours) $(3.10\pm 1.81)\times 10^{-7}$ $(1.09\pm 0.60)\times 10^{-8}$ $(1.11\pm 0.41)\times 10^{-9}$ $(1.13\pm 0.70)\times 10^{-9}$ $(1.14\pm 1.10)\times 10^{-9}$
EKAN-$\mathrm{O}(1,3)$ (Ours) $(3.14\pm 2.32)\times 10^{-7}$ $(1.46\pm 0.79)\times 10^{-8}$ $(1.50\pm 0.46)\times 10^{-9}$ $(1.50\pm 0.92)\times 10^{-9}$ $(1.54\pm 1.50)\times 10^{-9}$
Appendix G Implementation Details
G.1 Particle Scattering
@@
In particle scattering, we generate training sets of different sizes, and the corresponding test sets have the same sizes as the training sets. We train EKAN using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size <1000<1000, we set the number of epochs to $7000$, while for datasets with the training set size $\geq 1000$, we set the number of epochs to $15000$, which is sufficient for the MSE loss to converge to the minimum. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.
G.2 Three-Body Problem
@@
G.3 Top Quark Tagging
In top quark tagging, we train all models using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $\leq 1000$, we set the number of epochs to $1000$, while for datasets with the training set size >1000>1000, we set the number of epochs to $2000$, which is sufficient for the BCE loss to converge to the minimum. The grids of EKAN and KAN are updated every $5$ epochs and stop updating at the $50$th epoch. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.
Appendix H Additional Experiments
For top quark tagging, we increase the number of observed jet components for further comparison. We present the results using the $\mathrm{SO}^{+}(1,3)$-equivariant network as a representative, while the results of $\mathrm{SO}(1,3)$ and $\mathrm{O}(1,3)$-equivariant networks are very similar. We set the number of jet components to $n_{comp}=10$ and $n_{comp}=20$ respectively, and the experimental results are shown in Tables 8 and 9. Together with Table 3 in Section 6, we note that all models exhibit significant improvements in accuracy as the observed information increases. When $n_{comp}$ is larger, our EKAN achieves comparable results to the baselines, but it does not show superior performance. This suggests that EKAN may only have advantages in scenarios with insufficient observation information (as shown in Table 3), benefiting from its stronger generalization capability.
Table 8: Test accuracy ($\%$) of different models on the top quark tagging dataset ($n_{comp}=10$) with different training set sizes. We present the results in the format of mean $\pm$ std.
Models Training set size
$10^{2}$ $10^{2.5}$ $10^{3}$ $10^{3.5}$ $10^{4}$
EMLP-$\mathrm{SO}^{+}(1,3)$ $\mathbf{78.95\pm 2.24}$ $\mathbf{81.52\pm 0.48}$ $81.18\pm 0.42$ $82.50\pm 0.30$ $85.03\pm 0.04$
CGENN $71.82\pm 3.28$ $80.35\pm 0.57$ $79.85\pm 1.01$ $81.56\pm 0.23$ $84.17\pm 0.76$
EKAN-$\mathrm{SO}^{+}(1,3)$ (Ours) $78.57\pm 0.63$ $79.77\pm 0.78$ $\mathbf{82.90\pm 0.61}$ $\mathbf{84.83\pm 0.19}$ $\mathbf{87.14\pm 0.03}$
Table 9: Test accuracy ($\%$) of different models on the top quark tagging dataset ($n_{comp}=20$) with different training set sizes. We present the results in the format of mean $\pm$ std.
Models Training set size
$10^{2}$ $10^{2.5}$ $10^{3}$ $10^{3.5}$ $10^{4}$
EMLP-$\mathrm{SO}^{+}(1,3)$ $\mathbf{83.71\pm 0.49}$ $\mathbf{83.57\pm 0.69}$ $\mathbf{83.14\pm 0.35}$ $85.00\pm 0.35$ $86.81\pm 0.14$
CGENN $76.24\pm 1.28$ $82.34\pm 0.80$ $81.85\pm 0.40$ $84.67\pm 0.87$ $86.76\pm 0.50$
EKAN-$\mathrm{SO}^{+}(1,3)$ (Ours) $80.36\pm 1.99$ $81.25\pm 0.70$ $82.89\pm 1.18$ $\mathbf{86.21\pm 0.21}$ $\mathbf{89.30\pm 0.12}$